# Bochner’s technique in Einstein’s non-symmetric geometry

### v1

@@ Abstract
A. Einstein considered a manifold with a non-symmetric (0,2)-tensor $G$ and a linear connection $\nabla=(\Gamma^{k}_{ij})$ satisfying the condition $\dfrac{\partial G_{ij}}{\partial x^{k}}=\Gamma^{p}_{ik}G_{pj}+\Gamma^{p}_{kj}G_{ip}$. Guided by the construction of an almost Lie algebroid (on a vector bundle), we define the following concepts of Bochner’s technique for the Einstein’s non-symmetric geometry: the $\nabla^{f}$-connection, Bochner and Hodge $f$-Laplacians on tensors, the $f$-curvature operator and the Weitzenböck type curvature operator of Einstein’s connection. We prove Weitzenböck type decomposition formula and obtain vanishing results about the null space of the Bochner and Hodge $f$-Laplacians.
@@ 1 Introduction
Since 1923, A. Einstein worked on various versions of the Unified Field Theory (Non-symmetric Gravitational Theory – NGT), see [3]. This theory was intended to unite the gravitation theory, to which General Relativity (GR) is related, and the theory of electromagnetism. Introducing various versions of his NGT, A. Einstein used a complex basic tensor, with a symmetric real part and a skew-symmetric imaginary part. Beginning in 1950, A. Einstein used a real non-symmetric basic tensor $G=(G_{ij})$, whose symmetric part $g=(g_{ij})$ is associated with gravity, and the skew-symmetric part $F=(F_{ij})$ is associated with electromagnetism. The same is true for the symmetric part of the connection and torsion tensor, respectively. In his attempt to construct an unified field theory, (Einstein’s Gravitational Theory, briefly NGT), A. Einstein [3] considered a manifold $(M,G=g+F)$ with a linear connection $\nabla=(\Gamma^{k}_{ij})$ (not uniquely) determined by the condition $\dfrac{\partial G_{ij}}{\partial x^{k}}=\Gamma^{p}_{ik}G_{pj}+\Gamma^{p}_{kj}G_{ip}$ (called an Einstein’s connection), or, in coordinate-free form, see [4, 16]:
@@
A manifold $(M,G=g+F)$ equipped with an Einstein’s connection $\nabla$ different from the Levi-Civita connection $\nabla^{g}$ is called an NGT space, or Einstein’s non-symmetric geometry.
The idea of a non-symmetric basic tensor was later revisited by J.W. Moffat [6], allowing for richer gravitational dynamics. Results by T. Jansen and T. Prokopec [5] suggest potential improvements and challenges within NGT. M.I. Wanas et al. [15] studied path equations of a particle in Einstein’s non-symmetric geometry. Recent approaches to modified gravity often rely on differential geometry, incorporating torsion and non-metricity as natural extensions of GR framework. S. Ivanov and M. Zlatanović [4] investigated NGT using connections with skew-torsion and gave examples with the skew-symmetric part $F$ of $G$. V. Rovenski and M. Zlatanović [16] were the first to apply weak $f$-structures, in particular, weak almost contact structures, to NGT of skew-torsion with $F(X,Y)=g(X,fY)$. Weak metric structures (developed in recent works by V. Rovenski, see surveys [13, 14]) generalize the classical structures and are well suited for a study of $G=g+F$ in NGT with an arbitrary skew-symmetric tensor $F$.
@@
The Weitzenböck formula (2) allows us to extend the Hodge Laplacian to arbitrary tensors.
@@
In this paper, we continue our study [10, 11] and develop the Bochner technique for NGT spaces: the $f$-connection, $f$-divergence, $f$-curvature tensor, Bochner and Hodge $f$-Laplacians and Weitzenböck $f$-curvature operator (where $f$, see (5), is conjugate to $F$). Our results hold both in the case where $f$ is non-degenerate and in the case where $\ker f$ defines a singular distribution (i.e, $\dim\ker f>0$ at some points), which applies to the weak metric $f$-structure, see [13, 14].
The work has the Introduction and two sections, the References include 16 items. In Section 2, we recall the basics of Einstein’s non-symmetric geometry and the almost Lie algebroid structure. In Section 3 (consisting of five subsections), we define for a manifold $(M,G=g+F)$ with an Einstein’s connection $\nabla$, the derivatives $\nabla^{f}$ and $d^{f}$ (Section 3.1), the $f$-divergence and their $L^{2}$ adjoint operators on tensors (Section 3.2), and Bochner and Hodge $f$-Laplacians on tensors and differential forms (Section 3.3). In Section 3.4, making some assumptions about $f$, we define the $f$-curvature operator $R^{f}$ of $\nabla^{f}$, and in Section 3.5, we define the Weitzenböck type curvature operator on tensors, prove Weitzenböck type decomposition formula and obtain vanishing results.
@@ 2 Preliminaries
Here we recall the basics of Einstein’s NGT and the almost Lie algebroid structure.
@@ 2.1 Einstein’s non-symmetric geometry
The basic (0,2)-tensor $G=(G_{ij})$ in a “generalized” Riemannian manifold $(M,G)$ is, in general, non-symmetric. It decomposes in two parts, $G=g+F$, the symmetric part $g=(g_{ij})$ (called Riemannian metric) and the skew-symmetric part $F=(F_{ij})$ (called fundamental 2-form), where
@@
We assume that the symmetric part, $g$, is non-degenerate of arbitrary signature, and the skew-symmetric part, $F\neq 0$, has arbitrary (not necessarily constant) rank, e.g., is non-degenerate. Therefore, we obtain a (1,1)-tensor ${f}\neq 0$ (of not necessarily constant rank) determined by
@@
Since $F$ is skew-symmetric, the tensor f{f} is also skew-symmetric: g(fX,Y)=g(X,fY)g({f}X,Y)=-g(X,{f}Y).
We consider linear connections $\nabla$ on $M$ with a torsion (1,2)-tensor $T(X,Y)=\nabla_{X}Y-\nabla_{Y}X-[X,Y]$. We denote the torsion (0,3)-tensor with respect to $g$ by the same letter,
$T(X,Y,Z):=g(T(X,Y),Z).$ (6)
A linear connection on a manifold $(M,G=g+F)$ is completely determined by the torsion tensor and the covariant derivative $\nabla g$ of the symmetric part $g$ of $G$, see [4].
The covariant derivative an $(p,k)$-tensor $S$, where $p=0,1$, as an $(p,k+1)$-tensor $\nabla S$:
@@
Using (4), (5), (6) and (7), the condition (1) can be presented in the following form, see [4]:
$(\nabla G)(X,Y,Z)=-G(T(X,Y),Z)\ \Longleftrightarrow\ (\nabla(g+F))(X,Y,Z)=-T(X,Y,Z)-T(X,Y,{f}Z).$ (8)
Separating symmetric and skew-symmetric parts of (8) (with respect to $Y$ and $Z$), we express the covariant derivatives $\nabla g$ and $\nabla F$ in terms of torsion:
@@
The difference (1,2)-tensor $K$ of a linear connection $\nabla$ and the Levi-Civita connection $\nabla^{g}$ is
$K_{X}Y:=\nabla_{X}Y-\nabla^{g}_{X}Y.$ (10)
By (9), the difference tensor and torsion of an Einstein’s connection $\nabla$ are related as
$K(X,Y,Z)=\frac{1}{2}\,\big[T(X,Y,Z)+T(X,Z,{f}Y)+T(Y,Z,{f}X)\big],$ (11)
where $K(X,Y,Z):=g(K_{X}Y,Z)$. Therefore, for an Einstein’s connection $\nabla$ we have
$K_{X}Y-K_{Y}X=T(X,Y).$ (12)
@@
Linear connections with skew-torsion, T(X,Y,Z)=T(X,Z,Y)=T(Y,X,Z)T(X,Y,Z)=-T(X,Z,Y)=-T(Y,X,Z), have become popular due to the relations with mathematical physics. In this case, the number of preserved supersymmetries is equal to the number of parallel spinors with respect to such a connection. S. Ivanov S. and M.Lj. Zlatanović, see [4], presented conditions for the existence and uniqueness of the Einstein’s connection with skew-torsion on a manifold $(M,G=g+F)$ and gave its explicit expression. In this case, the torsion is determined by $dF$: $T(X,Y,Z)=-\frac{1}{3}dF(X,Y,Z)$, and the difference tensor is given by the following formula:
@@
Let $(M,f,Q,g)$ be a weak nearly Kähler manifold, i.e., (14) and $(\nabla^{g}_{X}f)X=0$ are true, considered as a generalized Riemannian manifold $(M,G=g+F)$. Suppose that $\nabla$ is an Einstein’s connection on $(M,f,Q,g)$ with skew-torsion. Then $\nabla$ satisfies the $f$-torsion condition: $T(fX,Y)=T(X,fY)$, and torsion of $\nabla$ is determined by $dF$ as $T(X,Y,Z){=}-\frac{1}{3}dF(X,Y,Z)$, see Example 3.4 in [16]. In this case, (13) reduces to $K_{X}Y=\frac{1}{2}\,T(X,Y)$, and our Einstein’s connection $\nabla$ is recovered by the equality $\nabla_{X}Y=\nabla^{g}_{X}Y+\frac{1}{2}\,T(X,Y)$.
@@
For an Einstein’s connection $\nabla$ on a manifold $(M,G=g+F)$, the tensor $K(X,Y,Z)$ cannot be totally symmetric, see Remark 2. Thus, it is interesting to study its symmetries separately.
The skew-symmetry condition of $K_{X}$ that is, K(X,Y,Z)=K(X,Z,Y)K(X,Y,Z)=-K(X,Z,Y), by (11), reduces to
@@
$E:=\sum\nolimits_{\,i}K_{e_{i}}e_{i},\quad E^{*}:=\sum\nolimits_{\,i}K^{*}_{e_{i}}e_{i},$ (17)
@@ 2.2 The almost Lie algebroid structure
Here, we briefly recall the construction of an almost Lie algebroid, following [9, 10]. Lie algebroids (and Lie groupoids) constitute an active field of research in differential geometry. Roughly speaking, an (almost) Lie algebroid is a structure, where one replaces the tangent bundle $TM$ of a manifold $M$ with a new smooth vector bundle $\pi_{\cal E}\colon{\cal E}\to M$ of rank $k$, i.e., with a fibre $\mathbb{R}^{k}$, over $M$ with similar properties. Many geometrical notions, which involve $TM$, were generalized to the context of Lie algebroids.
Definition 1.
@@
$\mathcal{J}_{\rho}(X,Y,Z)=\sum\nolimits_{\,\mathrm{cycl.}}[X,[Y,Z]_{\rho}\,]_{\rho},$
which prevents a bracket to satisfy the Jacobi identity, is called the Jacobiator of the bracket.
By (18), third formula, we get $\rho\,\mathcal{J}_{\rho}=0$. An almost Lie algebroid is a Lie algebroid provided that $\mathcal{J}_{\rho}=0$. Lie algebroids deal with integrable distributions (foliations). Almost Lie algebroids are closely related to singular distributions, i.e., given by $\rho({\cal E})$, see [10].
Note that axiom (18), third formula, is equivalent to vanishing of the following operator:
@@
+ $+\mathrel{\mathop{\sum}\nolimits_{\,0\leq i<j\leq k}}(-1)^{i+j}\omega([X_{i},X_{j}]_{\rho},X_{0},\ldots,X_{i-1},X_{i+1},\ldots,X_{j-1},X_{j+1},\ldots,X_{k}),$ (20)
@@
– a Lie algebroid if and only if $(d^{\rho})^{2}f=0$ and $(d^{\rho})^{2}\,\omega=0$ for $f\in C^{\infty}(M)$ and $\omega\in\Lambda^{1}(M)$.
Definition 2.
@@
The following equality holds:
$\sum\nolimits_{\,\mathrm{circ.}}R^{\rho}_{\,X,Y}Z=\sum\nolimits_{\,\mathrm{circ.}}[(\nabla^{\rho}_{X}\,T^{\rho})(Y,Z)+T^{\rho}(T^{\rho}(X,Y),Z)]+\mathcal{J}_{P}(X,Y,Z).$
We will apply the construction of an almost Lie algebroid to a manifold $(M,G=g+F)$ with an Einstein’s connection, namely, in the rest of paper, we assume ${\cal E}=TM$ and $\rho=f$, see (5).

### v2

@@ Abstract
A. Einstein considered a manifold with a non-symmetric (0,2)-tensor $G=g+F$, where $g$ is a Riemannian metric and $F\neq 0$, and a connection $\nabla$ with torsion $T$ such that $(\nabla_{X}G)(Y,Z)=-G(T(X,Y),Z)$. Guided by the almost Lie algebroid construction on a vector bundle, we define the basic concepts of Bochner’s technique for Einstein’s non-symmetric geometry, give a clear example of the Einstein’s connection $\nabla$, prove Weitzenböck type decomposition formula and obtain vanishing results about the null space of the Bochner and Hodge type Laplacians.
@@ 1 Introduction
Since 1923, A. Einstein worked on various versions of the Unified Field Theory (Non-symmetric Gravitational Theory – NGT), see [3]. This theory was intended to unite the gravitation theory, to which General Relativity (GR) is related, and the theory of electromagnetism. Introducing various versions of his NGT, A. Einstein used a complex basic tensor, with a symmetric real part and a skew-symmetric imaginary part. Beginning in 1950, A. Einstein used a real non-symmetric basic (0,2)-tensor $G$ whose symmetric part $g$ is associated with gravity, and the skew-symmetric part $F$ is associated with electromagnetism. The same is true for the symmetric part of the connection and torsion tensor, respectively. In his attempt to construct an unified field theory, (Einstein’s Gravitational Theory, briefly NGT), A. Einstein [3] considered a manifold $(M,G=g+F)$, where $g$ is a Riemannian metric and $F\neq 0$, with a linear connection $\nabla$ with torsion $T$ such that $(\nabla_{X}G)(Y,Z)=-G(T(X,Y),Z)$ (called an Einstein’s connection), or, equivalently, see [4, 16]:
@@
A manifold $(M,G=g+F)$ equipped with an Einstein’s connection $\nabla$ different from the Levi-Civita connection $\nabla^{g}$ is called an NGT space, or Einstein’s non-symmetric geometry. The idea of a non-symmetric basic tensor was later revisited by J.W. Moffat [6], allowing for richer gravitational dynamics. Results by T. Jansen and T. Prokopec [5] suggest potential improvements and challenges within NGT. M.I. Wanas et al. [15] studied path equations of a particle in Einstein’s non-symmetric geometry. Recent approaches to modified gravity often rely on differential geometry, incorporating torsion and non-metricity as natural extensions of GR framework. S. Ivanov and M. Zlatanović [4] investigated NGT using connections with totally skew-symmetric torsion and gave examples with the skew-symmetric part $F$ of $G$. V. Rovenski and M. Zlatanović [16] were the first to apply weak $f$-structures, in particular, weak almost contact structures, to NGT of totally skew-symmetric torsion with $F(X,Y)=g(X,fY)$. Weak metric structures (developed in recent works by V. Rovenski, see surveys [13, 14]) generalize the classical structures and are well suited for a study of $G=g+F$ in NGT with an arbitrary skew-symmetric tensor $F$.
@@
The formula (2) allows us to extend the Hodge Laplacian to arbitrary tensors.
@@
In the paper, we consider Einstein’s connections, see (1), satisfying the additional condition $\nabla g=0$, hence the properties (ii) and (iii) of Lemma 1 below hold. We develop the Bochner technique for NGT spaces: the $f$-connection, $f$-divergence, $f$-curvature tensor, $f$-Laplacians and Weitzenböck $f$-curvature operator (where $f$, see (5), is conjugate to $F$). Our results hold both in the case where $f$ is non-degenerate and in the case where $\ker f$ defines a singular distribution (i.e, $\dim\ker f>0$ at some points), which applies to the weak metric $f$-structure, see [13, 14].
The work has the Introduction and two sections, the References include 16 items. In Section 2, we recall the basics of Einstein’s non-symmetric geometry and the almost Lie algebroid structure. Section 3 consists of five subsections. In Section 3.1, for a manifold $(M,G=g+F)$ with an Einstein’s connection $\nabla$, we define the derivative $\nabla^{f}$, the $f$-bracket and give a clear example of the Einstein’s connection $\nabla$. In Section 3.2, we define the skew-symmetric $f$-divergence and their $L^{2}$ adjoint operators on tensors. In Section 3.3, we introduce Bochner and Hodge $f$-Laplacians on tensors and differential forms. In Sections 3.4 and 3.5, we define the $f$-curvature operator $R^{f}$ of $\nabla^{f}$, the Weitzenböck type curvature operator on tensors, prove Weitzenböck type decomposition formula and obtain vanishing results.
@@ 2 Preliminaries
Here, we discuss Einstein’s non-symmetric geometry and the almost Lie algebroid structure.
@@ 2.1 Einstein’s non-symmetric geometry
The basic (0,2)-tensor $G$ in a “generalized” Riemannian manifold $(M,G)$ is non-symmetric. It decomposes in two parts, $G=g+F$, the symmetric part $g>0$ (called Riemannian metric) and the skew-symmetric part $F\neq 0$ (called fundamental 2-form), where
@@
Therefore, we obtain a (1,1)-tensor ${f}\neq 0$ (of not necessarily constant rank) determined by
@@
Since $F$ is skew-symmetric, the tensor f{f} is also skew-symmetric: g(fX,Y)=g(X,fY)g({f}X,Y)=-g(X,{f}Y). We consider linear connections $\nabla$ on $M$ with a torsion (1,2)-tensor
$T(X,Y)=\nabla_{X}Y-\nabla_{Y}X-[X,Y].$
Denote the torsion (0,3)-tensor with respect to $g$ by the same letter, $T(X,Y,Z):=g(T(X,Y),Z)$.
The covariant derivative an $(p,k)$-tensor $S$ for $p=0,1$, as an $(p,k+1)$-tensor $\nabla S$, see [8]:
@@
Using (4), (5) and (6), we present the condition (1) in the following form, see [4]:
@@
The above equations yield, see [4, Eq. (3.4)],
$(\nabla\,g)(Z,X,Y)=(\nabla\,F)(X,Y,Z)+(\nabla\,F)(Y,X,Z).$ (7)
Lemma 1.
For an Einstein’s connection $\nabla$, the following conditions are equivalent:
$(i)$ $\nabla g=0$,
$(ii)$ K(X,Y,Z)=K(X,Z,Y)K(X,Y,Z)=-K(X,Z,Y),
$(iii)$ $(\nabla_{X}F)(Y,Z)=-(\nabla_{Y}F)(X,Z)$, or $(\nabla_{X}f)Y=-(\nabla_{Y}f)X$.
Proof.
We calculate using (6) and (8),
$(\nabla_{X}\,g)(Y,Z)$ $=X(g(Y,Z))-g(\nabla^{g}_{X}Y+K_{X}Y,Z)-g(Y,\nabla^{g}_{X}Z+K_{X}Z)$
$=(\nabla^{g}_{X}\,g)(Y,Z)-g(K_{X}Y,Z)-g(K_{X}Z,Y).$
By this and $\nabla^{g}\,g=0$, the conditions (i) and (ii) are equivalent. By the equality (7), the conditions (i) and (iii) are equivalent. ∎
The difference tensor $K$ of a linear connection $\nabla$ and $\nabla^{g}$ is given by
$K_{X}Y:=\nabla_{X}Y-\nabla^{g}_{X}Y.$ (8)
It is called the contorsion tensor when $\nabla$ preserves the metric tensor, $\nabla g=0$, in Riemann-Cartan geometry. Denote the torsion (0,3)-tensor with respect to $g$ by the same letter, $K(X,Y,Z):=g(K_{X}Y,Z)$. In the case $\nabla g=0$, the tensors $K$ and $T$ are expressed linearly through each other:
$T(X,Y)$ $=K_{X}Y-K_{Y}X,$ (9)
$2\,K(Y,Z,X)$ $=T(X,Y,Z)+T(Y,Z,X)+T(X,Z,Y).$ (10)
@@
By (11), the tensor $K$ and the torsion $T$ of an Einstein’s connection $\nabla$ are related as
2K(X,Y,Z)=T(X,Y,Z)+T(X,Z,fY)+T(Y,Z,fX).2\,K(X,Y,Z)=T(X,Y,Z)+T(X,Z,{f}Y)+T(Y,Z,{f}X). (12)
Comparing (10) and (12) yields the condition for the torsion of an Einstein’s connection
$T(Z,X,Y)+T(Z,Y,X)=T(X,Z,{f}Y)+T(Y,Z,{f}X).$
@@
Linear connections with totally skew-symmetric torsion, T(X,Y,Z)=T(X,Z,Y)=T(Y,X,Z)T(X,Y,Z)=-T(X,Z,Y)=-T(Y,X,Z), have become popular due to the relations with mathematical physics. In this case, the number of preserved supersymmetries is equal to the number of parallel spinors with respect to such a connection. S. Ivanov S. and M.Lj. Zlatanović, see [4], presented conditions for the existence and uniqueness of the Einstein’s connection with totally skew-symmetric torsion on a manifold $(M,G=g+F)$ and gave its explicit expression. In this case, the torsion is determined by $dF$: $T(X,Y,Z)=-\frac{1}{3}dF(X,Y,Z)$, and the difference tensor is given by the following formula:
@@
Let $(M,f,Q,g)$ be a weak nearly Kähler manifold, i.e., (14) and $(\nabla^{g}_{X}f)X=0$ are true, considered as a generalized Riemannian manifold $(M,G=g+F)$. Suppose that $\nabla$ is an Einstein’s connection on $(M,f,Q,g)$ with totally skew-symmetric torsion. Then $\nabla$ satisfies the $f$-torsion condition: $T(fX,Y)=T(X,fY)$, and torsion of $\nabla$ is determined by $dF$ as $T(X,Y,Z){=}-\frac{1}{3}dF(X,Y,Z)$, see Example 3.4 in [16]. In this case, (13) reduces to $K_{X}Y=\frac{1}{2}\,T(X,Y)$, and our Einstein’s connection $\nabla$ is recovered by the equality $\nabla_{X}Y=\nabla^{g}_{X}Y+\frac{1}{2}\,T(X,Y)$.
@@
For an Einstein’s connection on a manifold $(M,G=g+F)$, the tensor $K(X,Y,Z)$ is not totally symmetric, see Remark 2. Thus, it is interesting to study its particular symmetries. The skew-symmetry of $K_{X}$ i.e., K(X,Y,Z)=K(X,Z,Y)K(X,Y,Z)=-K(X,Z,Y), by (12), reduces to
@@
Example 2.
If a manifold $(M,G=g+F)$ admits an Einstein’s connection $\nabla$ with totally skew-symmetric torsion, and the $f$-torsion condition holds: $T(fX,Y)=T(X,fY)$, see Example 1, then (16) is true. Indeed, for such a $\nabla$ the $f$-torsion condition implies
T(fY,Z)=T(Y,fZ)=fT(Y,Z).T(fY,Z)=T(Y,fZ)=-f\,T(Y,Z). (17)
Using (17) and the totally skew-symmetric torsion property in the LHS of (15) gives zero.
@@
$E:=\sum\nolimits_{\,i}K_{e_{i}}e_{i},\quad E^{*}:=\sum\nolimits_{\,i}K^{*}_{e_{i}}e_{i},$
@@ 2.2 The almost Lie algebroid structure
Here, we briefly recall the construction of an almost Lie algebroid, following [9, 10]. Lie algebroids constitute an active field of research in differential geometry. Roughly speaking, an (almost) Lie algebroid is a structure, where one replaces the tangent bundle $TM$ of a manifold $M$ with a new smooth vector bundle $\pi_{\cal E}\colon{\cal E}\to M$ of rank $k$, i.e., with a fibre $\mathbb{R}^{k}$, over $M$ with similar properties. Many geometrical notions, which involve $TM$, were generalized to the context of Lie algebroids.
@@
+ $+\sum\nolimits_{\,0\leq i<j\leq k}(-1)^{i+j}\omega([X_{i},X_{j}]_{\rho},X_{0},\ldots,X_{i-1},X_{i+1},\ldots,X_{j-1},X_{j+1},\ldots,X_{k}),$ (20)
@@
– a Lie algebroid if and only if $(d^{\rho})^{2}f=0=(d^{\rho})^{2}\,\omega$ for $f\in C^{\infty}(M)$ and $\omega\in\Lambda^{1}(M)$.
@@
We will apply the construction of an almost Lie algebroid to a manifold $(M,G=g+F)$ with an Einstein’s connection, namely, we assume ${\cal E}=TM$ and $\rho=f$, see (5).